A. A hotels receives a room reservation that indicates that the guest is making the reservation will pay $200 for their one night stay. The pays a 15% travel agent commission for the reservation, and the hotel's franchisor will assess total fees of 12% for the reservation. What is the hotel's net ADR yield for this one night reservation?

A. 77%
B. 75%
C. 79%
D. 73%

B. In which scenario would a forecast most likely be inaccurate because of its intended purpose?

A. When it has been recently modified to reflect actual performance.
B. When the time period it addresses is in the very near future.
C. When it is developed to make management look good to a business' owners.
D. When it is developed based primarily on historical data.

C. A manager is considering staying open one extra hour per day and wants to calculate the minimum sales point (MSP). The manager's variable costs are 40%, food costs are 30%, and fixed costs are 60%. The minimum labor cost for staying open one extra hour is $125.00. Rounded up to the nearest dollar, what is the manager's MSP?

D. What is the mathematical relationship between selling price, food cost %, and contribution margin?

A. Selling price x (1.00 - food cost %) = contribution margin
B. Selling price / (1.00 + food cost %) = contribution margin
C. Selling price x (1.00 + food cost %) = contribution margin
D. Selling price / (1 - food cost %) = contribution margin
A. $407
B. $457
C. $417
D. $437

E. A restaurant manager achieved $72,500 in revenue last January. For this coming January, they predict a 5.5% increase in monthly sales. What will be this manager's revenue forecast for this coming January?

A. $74,488
B. $78,488
C. $76,488
D. $72,488

F. A hotel's F&B department will host a banquet for 100 guests. The guests are to be charged $75 per person for food and $35 per person for alcoholic beverages. The service charge for the event will be 20%. What is the amount of food and beverage income (revenue) that will result from hosting this event?

A. $8,400
B. $13,200
C. $11,000
D. $16,300
Answer #1

Please upvote !!!

1.

A hotels receives a room reservation that indicates that the guest is making the reservation will pay $200 for their one night stay. The pays a 15% travel agent commission for the reservation, and the hotel's franchisor will assess total fees of 12% for the reservation. What is the hotel's net ADR yield for this one night reservation?

Ans -

Franchisr Fees = 12 %

Travel Agent Fees - 15%

Net Fees = 12+15 = 27%

So, the hotels net ADR = 100 - 27% = 73%

2

In which scenario would a forecast most likely be inaccurate because of its intended purpose?

Ans - C. When it is developed to make management look good to a business' owners. - Because such kind of forecasts are rife with inaccuracies and subjective judgment.

3.

A manager is considering staying open one extra hour per day and wants to calculate the minimum sales point (MSP). The manager's variable costs are 40%, food costs are 30%, and fixed costs are 60%. The minimum labor cost for staying open one extra hour is $125.00. Rounded up to the nearest dollar, what is the manager's MSP?

Ans :-

Fixed Cost = 60%

Variable Cost = 40% --( Food cost/ labour cost = 30% and Msc = 10%)

Labour Cost = $125 / hour

So, min MSP = 125/0.3 = $ 417

4. What is the mathematical relationship between selling price, food cost %, and contribution margin?

Answer - Selling price x (1.00 - food cost %) = contribution margin

5 A restaurant manager achieved $72,500 in revenue last January. For this coming January, they predict a 5.5% increase in monthly sales. What will be this manager's revenue forecast for this coming January?

Ans - $72,500 * 1.055 = $76487.5 -- Option C.

F. A hotel's F&B department will host a banquet for 100 guests. The guests are to be charged $75 per person for food and $35 per person for alcoholic beverages. The service charge for the event will be 20%. What is the amount of food and beverage income (revenue) that will result from hosting this event?

Ans :- Total Revenue per user = 75+35 = % 110

Revenue after commission = 1.2* 110 = $ 132

Number of guests = 100

Hence , revenue from hosting the event =$ 13200 --- Option B

PS- Please upvote if you find the answer useful and comment in case of any doubts. Request you to kindly leave a comment in case of any concerns as it would help refine the solution as opposed to a downvote.